# Model Tensor Planning

### v1

@@ A.2 Exploration Versus Exploitation Discussion
We investigate the tensor sampling Definition 1 (with linear interpolation) versus MPPI sampling with horizon $T$, corresponding to global exploration versus local exploitation behaviors from the current system state. In particular, we remark on the entropy of path distributions in both methods in the discretized control setting with equal time intervals. Further investigation on the continuous control setting is left for future work.
@@
Consider ${\mathcal{G}}(M,N)$, each node in layer $L_{i}$ is sampled independently from a uniform distribution over ${\mathbb{U}}$, and path candidates are equivalent to sequences of node indices ${\bm{\tau}}\sim\tau=(i_{1},i_{2},\dots,i_{M})\in\{1,\dots,N\}^{M}$ (cf. Algorithm 1). Let ${\mathcal{S}}$ denote the set of all index sequences representing valid paths through the graph. The uniform distribution over ${\mathcal{S}}$ is given by $P_{{\mathcal{G}}}({\bm{\tau}})=1/|{\mathcal{S}}|,\text{ where }|{\mathcal{S}}|=N^{M}$. Hence, the entropy of tensor sampling is
@@
Indeed, as $M,N\to\infty$, the entropy $H(P_{{\mathcal{G}}})\to\infty$, and the distribution over sampled paths in ${\mathcal{G}}$ becomes maximum entropy over ${\mathbb{F}}_{T}$ among all discrete path distributions. Theorem 1 implies that ${\mathbb{F}}_{T}\to{\mathbb{F}}$ as $M,N\to\infty$, and thus tensor sampling distribution becomes maximum entropy over ${\mathbb{F}}$.
@@ A.3 Task Details
A.4 Experiment Details
Comparison Experiment. We summarize the simulation settings for comparison experiments (cf. Section 4.2) in Table 2, and the hyperparameters used for MTP variants in Table 4. Tables 4 summarize the hyperparameters used for MPPI and PS (temperature is not relevant for PS). The same noise $\sigma$ is used for MTP local samples. For evolutionary strategies (DE and OpenAI-ES), we use default hyperparameters in evosax (Lange, 2023).
@@ A.4 Experiment Details
Design Ablation. We conducted the sweep on $\beta\in[0,1]$ and the number of elites to experimentally study the algorithmic design. The MTP hyperparameters in Section 4.3 are the same as in Table 4. Each setting was evaluated with $4$ random seeds.
Sensitivity Ablation. We conducted $\beta$-mixing rate ablation study (cf. Section 4.4) by varying the $\beta$ across different MTP variants. The MTP hyperparameters in Section 4.4 are the same as in Table 4. Each configuration was evaluated with $4$ random seeds.
@@ A.5 Additional Ablation & Performance Benchmarks
B-spline Degree Ablation. We investigate the sensitivity of MTP performance over B-spline interpolation degrees. The MTP hyperparameters are the same as in Table 4. Each setting was evaluated using $5$ random seeds.
@@
Figure 7: MTP-Bspline degree ablation. In G1-Walk, the Unitree G1 controlled Bspline degrees all roughly fall at $500$ time steps.
@@
In Fig. 7, we investigate the sensitivity of MTP performance over B-spline interpolation degrees. Results consistently show minimal performance differences across degrees ($p=2$ to $p=5$) in terms of both cost and velocity magnitude curves across different tasks. Given this insensitivity, we select the lower computational complexity B-spline $p=2$ as our default choice for the MTP-Bspline method.
@@
Sweep $M,N$ Ablation. We conducted an additional ablation to study the effect of varying $M$ and $N$ values of MTP variants on the Navigation task. We set $\beta=1$ to use full tensor sampling. Each configuration was evaluated with $4$ random seeds.
@@
Figure 8: Sweep $M,N$ on Navigation environment with $B=256$ to investigate the interplay between number of batch sample $B$, number of layer $M$, and number of control-waypoints per layer $N$. Each data point is the success rate over $4$ seeds. The environment setting is as in Section A.3.
@@
According to Fig. 8, there exists a sweet point in selecting the number of layers $M$ in tensor sampling. Roughly, for all MTP variants, increasing $M$ initially improves task performance, as more layers provide sufficient path complexity, allowing the planner to escape local minima and generate diverse, globally exploratory trajectories. However, beyond a certain point, further increasing $M$ degrades performance. This is due to the exponential growth in the number of possible paths ${\mathcal{O}}(N^{M})$, while the rollout budget $B$ remains fixed. As a result, the sampled trajectory density becomes sparse relative to the vast number of paths, reducing effective coverage and leading to diminished exploration and performance. On the other hand, increasing $N$ consistently improves performance by densifying the search at each layer. However, this comes at a higher computational cost. Therefore, a careful balance must be struck between $M$ and $N$ to maintain real-time control and effective control exploration.
@@
Planning Performance. Table 5 benchmarks the performance of our JAX-based implementation by measuring the wall-clock time of the JIT-compiled planning function on G1-Standup task. This function includes sampling, trajectory rollout, cost evaluation, and parameter updates. The task setting is similar to Table 2, but we set the sim step per plan to $1$ with $B=128$. The results show that the initial JIT compilation incurs a significant one-time cost, as expected for MuJoCo XLA pipelines. However, after compilation, the per-step planning rates across MTP, MPPI, PS, and evolutionary baselines are roughly similar.
@@
Table 5: JAX implementation benchmark on G1-Standup, evaluated with $5$ seeds on an Nvidia RTX 3090.
@@
JIT Time [s] $76.4\pm 1.2$ $74.62\pm 2.5$ $72.35\pm 4.5$ $73.87\pm 4.2$ $69.87\pm 1.2$ $73.62\pm 3.1$
Planning Time [ms] $2.7\pm 0.3$ $2.7\pm 0.4$ $3.1\pm 0.7$ $2.6\pm 0.2$ $2.9\pm 0.5$ $3.2\pm 0.6$
@@
Softmax Update Effect. Fig. 9 illustrates the impact of applying softmax weighting on elite candidates for updating the mean and standard deviation of control trajectories. The left plot demonstrates smoother and lower-variance control updates over time compared to updates without softmax weighting shown on the right. The smooth and stable updates afforded by softmax weighting are essential when effectively mixing global and local samples, highlighting its critical role in the MTP performance.
Figure 9: Softmax weighting ablation on PushT environment. Both control update trajectories converge to near-zero means with large variance at $100$ timesteps, signifying task completion.

### v2

@@ A.2 Exploration Versus Exploitation Discussion
We investigate the tensor sampling Definition˜1 (with linear interpolation) versus MPPI sampling with horizon $T$, corresponding to global exploration versus local exploitation behaviors from the current system state. In particular, we remark on the entropy of path distributions in both methods in the discretized control setting with equal time intervals. Further investigation on the continuous control setting is left for future work.
@@
Consider ${\mathcal{G}}(M,N)$, each node in layer $L_{i}$ is sampled independently from a uniform distribution over ${\mathbb{U}}$, and path candidates are equivalent to sequences of node indices ${\bm{\tau}}\sim\tau=(i_{1},i_{2},\dots,i_{M})\in\{1,\dots,N\}^{M}$ (cf. Algorithm˜1). Let ${\mathcal{S}}$ denote the set of all index sequences representing valid paths through the graph. The uniform distribution over ${\mathcal{S}}$ is given by $P_{{\mathcal{G}}}({\bm{\tau}})=1/|{\mathcal{S}}|,\text{ where }|{\mathcal{S}}|=N^{M}$. Hence, the entropy of tensor sampling is
@@
Indeed, as $M,N\to\infty$, the entropy $H(P_{{\mathcal{G}}})\to\infty$, and the distribution over sampled paths in ${\mathcal{G}}$ becomes maximum entropy over ${\mathbb{F}}_{T}$ among all discrete path distributions. Theorem˜1 implies that ${\mathbb{F}}_{T}\to{\mathbb{F}}$ as $M,N\to\infty$, and thus tensor sampling distribution becomes maximum entropy over ${\mathbb{F}}$.
@@ A.3 Task Details
A.4 Experiment Details & Discussions
Comparison Experiment. We summarize the simulation settings for comparison experiments (cf. Section˜4.2) in Table˜2, and the hyperparameters used for MTP variants in Table 4. Tables 4 summarize the hyperparameters used for MPPI and PS (temperature is not relevant for PS). The same noise $\sigma$ is used for MTP local samples. For evolutionary strategies (DE and OpenAI-ES), we use default hyperparameters in evosax (Lange, 2023).
@@ A.4 Experiment Details & Discussions
Design Ablation. We conducted the sweep on $\beta\in[0,1]$ and the number of elites to experimentally study the algorithmic design. The MTP hyperparameters in Section˜4.3 are the same as in Table˜4. Each setting was evaluated with 444 random seeds.
Sensitivity Ablation. We conducted $\beta$-mixing rate ablation study (cf. Section˜4.4) by varying the $\beta$ across different MTP variants. The MTP hyperparameters in Section˜4.4 are the same as in Table˜4. Each configuration was evaluated with 444 random seeds.
Experimental Discussions. In tasks with well-shaped or dense reward structures and moderately nonlinear dynamics, exploration becomes less critical and nominal sampling methods (e.g., MPPI, PS) often suffice—explaining MPPI’s strong performance in G1-Standup, which benefits from fully actuated dynamics and informative rewards. However, in more challenging scenarios involving sparse rewards or highly nonlinear dynamics, such as PushT and under domain shifts in Crane, these locally guided strategies tend to struggle with inadequate exploration, often converging to suboptimal solutions or showing high performance variance. A representative case is the Navigation task (Figure 3), where the agent must discover velocity sequences to bypass obstacles and reach the goal—a setting in which local Gaussian sampling clearly fails by getting trapped in local minima. To address these challenges, MTP introduces a structured high-entropy sampling mechanism along with a simple yet effective $\beta$-mixing strategy that balances global exploration and local exploitation. With careful tuning of $\beta$, $M$, and $N$, MTP demonstrates robust and consistent performance across diverse tasks.
Mixing Rate Tuning. $\beta$ determines the ratio between exploratory (tensor sampling) and exploitative (nominal sampling) samples, which is delicate to tune. As shown in our ablations in Fig.˜6 and Fig.˜5, high $\beta$ values (e.g., 0.50.50.5) introduce strong exploration, which may benefit tasks with sparse rewards and not requiring delicate fixed-point stability (e.g., PushT, Cube-In-Hand, Navigation). Conversely, in tasks requiring high stability or precise actuation, such as G1-Standup or G1-Walk, lower $\beta$ values (e.g., 0.050.050.05-0.10.10.1) tend to yield better performance by favoring consistent behavior while still injecting enough exploration to escape suboptimal solutions.
@@ A.5 Additional Ablation & Performance Benchmarks
B-spline Degree Ablation. We investigate the sensitivity of MTP performance over B-spline interpolation degrees. The MTP hyperparameters are the same as in Table˜4. Each setting was evaluated using 444 random seeds.
@@
Figure 7: MTP-Bspline degree ablation. In G1-Walk, the Unitree G1 controlled Bspline degrees all roughly fall at 500500500 time steps.
@@
In Fig.˜7, we investigate the sensitivity of MTP performance over B-spline interpolation degrees. Results consistently show minimal performance differences across degrees ($p=2$ to $p=5$) in terms of both cost and velocity magnitude curves across different tasks. Given this insensitivity, we select the lower computational complexity B-spline $p=2$ as our default choice for the MTP-Bspline method.
@@
Sweep $M,N$ Ablation. We conducted an additional ablation to study the effect of varying $M$ and $N$ values of MTP variants on the Navigation task. We set $\beta=1$ to use full tensor sampling. Each configuration was evaluated with 444 random seeds.
@@
Figure 8: Sweep $M,N$ on Navigation environment with $B=256$ to investigate the interplay between number of batch sample $B$, number of layer $M$, and number of control-waypoints per layer $N$. Each data point is the success rate over 444 seeds. The environment setting is as in Section˜A.3.
@@
According to Fig.˜8, there exists a sweet point in selecting the number of layers $M$ in tensor sampling. Roughly, for all MTP variants, increasing $M$ initially improves task performance, as more layers provide sufficient path complexity, allowing the planner to escape local minima and generate diverse, globally exploratory trajectories. However, beyond a certain point, further increasing $M$ degrades performance. This is due to the exponential growth in the number of possible paths ${\mathcal{O}}(N^{M})$, while the rollout budget $B$ remains fixed. As a result, the sampled trajectory density becomes sparse relative to the vast number of paths, reducing effective coverage and leading to diminished exploration and performance. On the other hand, increasing $N$ consistently improves performance by densifying the search at each layer. However, this comes at a higher computational cost. Therefore, a careful balance must be struck between $M$ and $N$ to maintain real-time control and effective control exploration.
@@
Planning Performance. Table 5 benchmarks the performance of our JAX-based implementation by measuring the wall-clock time of the JIT-compiled planning function on G1-Standup task. This function includes a single sampling, single trajectory rollout, single cost evaluation, and single parameter update. Note that this benchmark only serves as an exemplary simulated planning function performance with JAX JIT. In practice, we might have multiple search refinements, or multiple parameter updates in the control loops. The task setting is similar to Table˜2, but we set the sim step per plan to 111 with $B=128$. The results show that the initial JIT compilation incurs a significant one-time cost, as expected for MuJoCo XLA pipelines. However, after compilation, the per-step planning rates across MTP, MPPI, PS, and evolutionary baselines are roughly similar with the same batch sample $B$, as also reflected in Table˜6, Table˜7, and Table˜8. The results confirm that the JIT [s] and Planning Time [ms] are algorithm-agnostic, which depends slightly on batch size $B$ (i.e., Planning Time [ms] logarithmically increases with $B$) and on the environment dynamics. All algorithms remain real-time feasible on GPU-accelerated hardware, when implemented with JAX and MuJoCo XLA.
@@
Table 5: JAX implementation benchmark on G1-Standup, evaluated with 555 seeds on an Nvidia RTX 3090.
@@
JIT Time [s] 76.4±1.276.4$\pm$ 1.276.4 ± 1.2 74.62±2.574.62$\pm$ 2.574.62 ± 2.5 72.35±4.572.35$\pm$ 4.572.35 ± 4.5 73.87±4.273.87$\pm$ 4.273.87 ± 4.2 69.87±1.269.87$\pm$ 1.269.87 ± 1.2 73.62±3.173.62$\pm$ 3.173.62 ± 3.1
Planning Time [ms] 2.7±0.32.7$\pm$ 0.32.7 ± 0.3 2.7±0.42.7$\pm$ 0.42.7 ± 0.4 3.1±0.73.1$\pm$ 0.73.1 ± 0.7 2.6±0.22.6$\pm$ 0.22.6 ± 0.2 2.9±0.52.9$\pm$ 0.52.9 ± 0.5 3.2±0.63.2$\pm$ 0.63.2 ± 0.6
Table 6: Planning performance of MTP-Akima. Averaged over 5 seeds on an Nvidia RTX 4090.
Batch Size $B$ JIT Time [s] Planning Time [ms]
PushT Crane Cube-In-Hand G1-Walk PushT Crane Cube-In-Hand G1-Walk
646464 15.815.815.8 32.532.532.5 38.238.238.2 65.365.365.3 1.7±0.11.7$\pm$ 0.11.7 ± 0.1 1.8±0.11.8$\pm$ 0.11.8 ± 0.1 8.1±0.38.1$\pm$ 0.38.1 ± 0.3 1.4±0.11.4$\pm$ 0.11.4 ± 0.1
128128128 12.712.712.7 31.531.531.5 36.636.636.6 58.558.558.5 1.6±0.11.6$\pm$ 0.11.6 ± 0.1 1.9±0.11.9$\pm$ 0.11.9 ± 0.1 10.2±0.710.2$\pm$ 0.710.2 ± 0.7 1.5±0.11.5$\pm$ 0.11.5 ± 0.1
256256256 16.316.316.3 31.431.431.4 38.838.838.8 57.157.157.1 2.0±0.22.0$\pm$ 0.22.0 ± 0.2 2.0±0.42.0$\pm$ 0.42.0 ± 0.4 14.7±0.814.7$\pm$ 0.814.7 ± 0.8 1.5±0.11.5$\pm$ 0.11.5 ± 0.1
Table 7: Planning performance of MPPI. Averaged over 5 seeds on an Nvidia RTX 4090.
Batch Size $B$ JIT Time [s] Planning Time [ms]
PushT Crane Cube-In-Hand G1-Walk PushT Crane Cube-In-Hand G1-Walk
646464 14.914.914.9 32.332.332.3 37.737.737.7 62.962.962.9 1.7±0.11.7$\pm$ 0.11.7 ± 0.1 1.7±0.11.7$\pm$ 0.11.7 ± 0.1 8.2±0.38.2$\pm$ 0.38.2 ± 0.3 1.4±0.11.4$\pm$ 0.11.4 ± 0.1
128128128 18.418.418.4 29.929.929.9 36.736.736.7 58.258.258.2 1.7±0.11.7$\pm$ 0.11.7 ± 0.1 1.8±0.11.8$\pm$ 0.11.8 ± 0.1 10.4±0.510.4$\pm$ 0.510.4 ± 0.5 1.4±0.11.4$\pm$ 0.11.4 ± 0.1
256256256 14.914.914.9 30.330.330.3 37.537.537.5 55.755.755.7 1.9±0.11.9$\pm$ 0.11.9 ± 0.1 1.8±0.11.8$\pm$ 0.11.8 ± 0.1 15.2±0.915.2$\pm$ 0.915.2 ± 0.9 1.4±0.11.4$\pm$ 0.11.4 ± 0.1
Table 8: Planning performance of OpenAI-ES. Averaged over 5 seeds on an Nvidia RTX 4090.
Batch Size $B$ JIT Time [s] Planning Time [ms]
PushT Crane Cube-In-Hand G1-Walk PushT Crane Cube-In-Hand G1-Walk
646464 15.115.115.1 32.132.132.1 39.239.239.2 61.561.561.5 1.7±0.11.7$\pm$ 0.11.7 ± 0.1 1.8±0.11.8$\pm$ 0.11.8 ± 0.1 8.2±0.38.2$\pm$ 0.38.2 ± 0.3 1.6±0.11.6$\pm$ 0.11.6 ± 0.1
128128128 19.019.019.0 30.130.130.1 38.438.438.4 57.957.957.9 1.7±0.11.7$\pm$ 0.11.7 ± 0.1 1.9±0.11.9$\pm$ 0.11.9 ± 0.1 10.3±0.410.3$\pm$ 0.410.3 ± 0.4 1.5±0.11.5$\pm$ 0.11.5 ± 0.1
256256256 16.116.116.1 30.930.930.9 42.342.342.3 55.155.155.1 2.0±0.12.0$\pm$ 0.12.0 ± 0.1 1.8±0.11.8$\pm$ 0.11.8 ± 0.1 14.9±0.714.9$\pm$ 0.714.9 ± 0.7 1.5±0.11.5$\pm$ 0.11.5 ± 0.1
@@
Softmax Update Effect. Fig.˜9 illustrates the impact of applying softmax weighting on elite candidates for updating the mean and standard deviation of control trajectories. The left plot demonstrates smoother and lower-variance control updates over time compared to updates without softmax weighting shown on the right. The smooth and stable updates afforded by softmax weighting are essential when effectively mixing global and local samples, highlighting its critical role in the MTP performance.
Figure 9: Softmax weighting ablation on PushT environment. Both control update trajectories converge to near-zero means with large variance at 100100100 timesteps, signifying task completion.